An aircraft carrier weighting 6 times 10^7 kg is t
Answer #1

Fluid Frictional Force acting on aircraft is F= \beta e^{\alpha v}

Now by newton's laws of motion, F= ma= mdv/dt

Comparing above two equations, ( as we neglected air friction)

mdv/dt= \beta e^{\alpha v}

dv/e^{\alpha v}= (\beta/m) dt

Integrating above equation, we get

  ln(e^{\alpha v})/ (ve^{\alpha v})= (\beta/m)t + C ( where C= constant)

  \alpha v/ (ve^{\alpha v})= (\beta/m)t + C  ( because: ln(e^x)= x)

  \alpha/e^{\alpha v}= (\beta /m)t+ C................(1)

Now given, at t=0s, v= 20 m/s

Using above values in equation,

\alpha/e^{20 \alpha }= 0+ C

or C= \alpha/e^{20 \alpha }

Putting above in equation1 ,

\alpha/e^{\alpha v}= (\beta /m)t + \alpha/e^{20 \alpha }

\alpha(1/e^{\alpha v}-1/e^{20 \alpha }) = (\beta /m)t

Now when v= 0, then value of time is

\alpha(1/e^{\alpha 0}-1/e^{20 \alpha }) = (\beta /m)t

\alpha(1/1-1/e^{20 \alpha }) = (\beta /m)t

Using given values, (0.1)(1/1-1/e^{20(0.1)) }) = ((3 \times 10^5) /(6 \times 10^7))t

  (0.1)(1-1/ 7.389}) = (5 \times 10^{-3})t

(0.1)(1-0.135) = (5 \times 10^{-3})t

t = 17.3 s ( ANS)